OpenStudy (anonymous):

simplify the given expression: square root of -6 (2+sqrt(-8) )??????

OpenStudy (anonymous):

What is sqrt(-8) ?

OpenStudy (anonymous):

\[ \sqrt{-6\left( 2+\sqrt{-8} \right)}\]

OpenStudy (anonymous):

is that it?

OpenStudy (anonymous):

sqrt means square root i dont no how to make the sign so i use the words

OpenStudy (anonymous):

That's fine, but I'm saying what is the value for the square root of -8

OpenStudy (anonymous):

no, the square root sign is over -6 and then in the parenthesses is (2+ the square root of -8)

OpenStudy (anonymous):

\[\sqrt{-6}(2+\sqrt{-8})\]

OpenStudy (anonymous):

That it?

OpenStudy (anonymous):

\[2+2isqrt{2}\] is your right side simplification

OpenStudy (anonymous):

Stupid radical didn't work

OpenStudy (anonymous):

... So are you gonna just keep posting all your homework questions and not work with someone to solve them or what?

OpenStudy (anonymous):

if you have a problem you can not reply

OpenStudy (anonymous):

simple as that

OpenStudy (anonymous):

Indeed. Or I can reply with unhelpful answers. Or I can complain.

OpenStudy (anonymous):

I have a wide array of possible options in front of me.

OpenStudy (anonymous):

o ok have fun idc waht you do